OpenStudy (anonymous):

Slopes of parallel and perpendicular lines: Problem type 2

OpenStudy (anonymous):

Consider the line y\[y=\frac{ 6 }{ 7 }x-6\] . (a) Find the equation of the line that is parallel to this line and passes through the point(6,-6) (b) Find the equation of the line that is perpendicular to this line and passes through the point (6,-6) .

OpenStudy (anonymous):

a.) parallel line means same slope so m=6/7 plug this and the point given to point-slope form y-(-6)=6/7(x-6) y+6=6/7x-36/7 y=6/7x-78/7 b.) slope is the negative inverse so m=-7/6 y+6=-7/6(x-6) y+6=-7/6x+7 y=-7/6x+1

OpenStudy (anonymous):

thank you :)